<e> Gerry needs a loan of 8000 € for his holiday trip. The loan period is 1 year and the payments are made quarterly. The loan has fixed principal payment and the interest rate is 3.8 % p.a. How much money Gerry needs for the third payment?


1
Expert's answer
2022-03-03T17:42:06-0500

"x=S(P+\frac{P}{(1+P)^N-1})"

S-value of loan

P- 1/100 of interest rate.

N - lian period in minth

P=3.8/(100x12)=0.00316666667

"x=8000(0.00316666667+\frac{0.00316666667}{(1+0.00316666667)^{12}-1})=680.47"